OpenStudy (anonymous):

After your sister turned 12 years old she began a rapid growth spurt during the next year of her life. After 3 months of growth her height was 1.5 meters and after 11 months of growth her height was 1.7 meters. Let's assume that your sister's height changes at a constant rate during her growth spurt. What was your sister's height in meters when her growth spurt began?

OpenStudy (perl):

so we can model an equation using (x,y) = (month after growth spurt, height) (3, 1.5) (11, 1.7) m = (1.7 - 1.5) / (11- 3) = .025 meters / month , is her growth rate

OpenStudy (perl):

so if 3 months after 12 years she is 1.5 meters, and each month she grow .025 meters, we can back track. 3 months before is 1.5 - 3 * .025 = 1.425 meters so she is 1.425 meters at 12 years old